Expenditure categories

Million in Birr

Personal Consumption Expenditure(C)

4500

Gross private domestic investment (I)

850

Gov. purchasing of goods and services (G)

1250

Export(X)

750

Import(M)

1000

Net Factor Income received from abroad (NFIRA)

950

Net Factor Income paid from abroad (NFIPA)

1050

Indirect Business tax (IBT)

450

Depreciation(D)

525


1
Expert's answer
2022-03-20T19:10:48-0400

a.) Gross Domestic Product (GDP)

"GDP = C + I + G + (X-M)"

"=4500+850 +1250+750-1000"

"=6350"


b.)Gross National Product (GNP)

"GNP=C+I+G+(X-M)+NFIR-NFIPA)"

"=4500+850 +1250+(750-1000)+ (950-1050)"

"=6250"


c.)C) Net National Income (NNP)

"NNP=GND-D"

"NNP=C + I + G + (X-M)-D"

"=4500+850 +1250+750-1000- 525"

"NNP=6350-525"

"= 5,825"


d.)National Income (NI)

"NI=NNP-IBT"

"NNP=6350-525-450"

"=5825-450"

"=53375"


e.) Trade Balance (NX)

"Nx=export -import"

"=750-1000"

"=-250"